3) There are 1,000 identical perfectly competitive real-estate firms selling office space in Syracuse, NY. The Marginal Cost of producing each square foot of space is constant and equal to $20. There are no fixed costs of production. So the firm’s short-run and long-run cost function is c(q) = 20q. The market demand is Q = 10,000 - 250p.

Suppose a monopolist decides to buy out all 1,000 firms at once and monopolize the real estate market in Syracuse. Assume he successfully buys everyone’s firm and that there is no cost advantage associated with being a monopolist. Hence the cost function for the monopolist ism C(Q) = 20Q.

i) What is the dead-weight loss due to this monopolization?
Answer #1

in perfect competition P= MC

TC= 20q

MC= 20

P= 20

Q= 10000 - 250(20)= 5000

MR= MC

Q= 10000 - 250p

250P = 10000 - Q

P= 40 - 0.004Q

TR= PQ = 40Q - 0.004Q^2

MR= 40 - 0.008Q

40 - 0.008Q = 20

Q= 2500

P= 40 - 0.004(2500)= 30

Deadweight loss = 0.5(30-20)(5000-2500) = 12500